### v2

Abstract
Multipartite quantum entanglement, as a core quantum resource, is fundamental to the advancement of quantum science and technology. In multipartite quantum systems, there are two kinds of quantum entanglement: $k$-nonseparability and $k$-partite entanglement. In this paper, we propose sufficient criteria for detecting $k$-nonseparability and $k$-partite entanglement by using the generalized Wigner-Yanase skew information and mutually unbiased measurements. Examples are given to demonstrate the detection capability and advantages of these criteria. As an application, an example of recognizing the networks by detecting the depth of quantum networks is given.
Multipartite quantum entanglement; $k$-separability; $k$-partite entanglement; generalized Wigner-Yanase skew information; mutually unbiased measurements
preprint: APS/123-QED
I Introduction
Quantum entanglement plays a crucial role in quantum information science and has important applications in many fields, such as quantum computing, quantum communication, quantum teleportation, quantum key distribution and quantum measurements 1 ; 2 ; 3 ; 4 ; 5 ; 6 . By virtue of entanglement, multi-qubits are allowed to establish intricate nonclassical correlations, which enhance the power of quantum computing, reinforce the security of quantum communication, and increase a remarkable measurement precision.
In quantum entanglement, entanglement detection is one of the main research topics. For bipartite quantum systems, many criteria for entanglement detection have been proposed (for example, see PPT ; GHB ; CW ; H ; bloch ; GGHE ; QH ; ZFFW ; ZZJW ; RL1 and the references therein). However, in multipartite quantum systems, the task of detecting entanglement grows increasingly complex and challenging with the expansion in the number of parties and the dimension of the systems. So far, there are two kinds of definitions for multipartite quantum entanglement. One is defined as “$k$-nonseparability” in terms of some subsystems bing entangled while others not, and another is defined as “$k$-partite entanglement” according to the difficulty of preparing entangled states, i.e., the depth of the entanglement GOTG ; DCT ; DC ; GYv .
For the detection of $k$-nonseparability, some methods are proposed, such as that based on semidefinite programming WCXMF , the entropic uncertainty relations BG , quantum Fisher information HLS ; HQGY , and the generalized Wigner-Yanase skew information HXGGY . For the detection of $k$-partite entanglement, several criteria have also been developed, such as that based on Wigner-Yanase skew information C and the permutation operators HGY . For other related results, the readers can refer to GHL ; LGY ; HPd ; WH ; RL ; VHE ; HQGY2 and the references therein. Although some approaches to entanglement detection have been developed from different viewpoints, it remains extremely difficult to identify $k$-nonseparability and $k$-partite entanglement for general mixed states.
In WF , the authors gave the concept of mutually unbiased measurements (MUMs) which offers a new perspective and tools for the flow of information and entanglement detection. MUMs have significant advantages in quantum information processing, quantum computing and quantum communication, including enhancing measurement accuracy, reducing errors, and improving entanglement detection efficiency. Based on MUMs, several criteria were proposed for multipartite quantum entanglement detection LGY ; CMF ; SC .
In this paper, we will continue to discuss the topic of detecting $k$-nonseparability and $k$-partite entanglement based on MUMs and the generalized Wigner-Yanase skew information. The structure of this paper is as follows. In Section 2, we give some background knowledge on several quantum informations, $k$-separability, and $k$-partite entanglement in multipartite quantum systems. Section 3 is devoted to derive criteria for detecting $k$-nonseparability and $k$-partite entanglement in multipartite quantum systems by combining MUMs and the generalized Wigner-Yanase skew information. In addition, we provide examples to demonstrate the effectiveness of these criteria as well as an application to distinct quantum networks. Section 4 is a conclusion.
II Preliminaries
In this section, we review some important concepts in quantum information, mutually unbiased measurements, $k$-separability and $k$-partite entanglement in multipartite quantum systems.
Throughout this paper, assume that $H_{i}$ is any complex Hilbert space with $\dim H_{i}=d_{i}<\infty$, $i=1,2,\cdots,N$, and let $H=H_{1}\otimes H_{2}\otimes\cdots\otimes H_{N}$. Denote by $\mathcal{B}(H)$ the algebra of all bounded linear operators and ${\mathcal{S}}(H)=\{\rho:\rho\geq 0,{\rm Tr}(\rho)=1\}$ is the closed convex set of all quantum states in $\mathcal{B}(H)$. $\mathbb{I}\in{\mathcal{B}}(H)$ and $\mathbb{I}_{i}\in{\mathcal{B}}(H_{i})$ stand for the identity operator on $H$ and $H_{i}$, respectively.
II.1 $k$-nonseparability and $k$-partite entanglement
For $2\leq k\leq N$, assume that $\gamma_{1},\cdots,\gamma_{k}$ are non-empty subsets of $\{1,2,\cdots,N\}$. We say that $\{\gamma_{1},\gamma_{2},\cdots,\gamma_{k}\}$ is a $k$-partition, denoted by $\gamma=\gamma_{1}|\gamma_{2}|\ldots|\gamma_{k}$, if it satisfies the conditions $\gamma_{i}\cap\gamma_{j}=\emptyset$ for $i\neq j$ and $\gamma_{1}\cup\cdots\cup\gamma_{k}=\{1,2,\cdots,N\}$.
A pure state $|\psi\rangle\langle\psi|\in\mathcal{S}(H)$ is said to be $k$-separable if there exists a $k$-partition $\gamma=\gamma_{1}|\gamma_{2}|\cdots|\gamma_{k}$ of $\{1,2,\cdots,N\}$ such that
$|\psi\rangle=\otimes_{i=1}^{k}|\psi_{\gamma_{i}}\rangle,$
where $|\psi_{i}^{\gamma_{i}}\rangle\langle\psi_{i}^{\gamma_{i}}|\in\mathcal{S}(H_{\gamma_{i}})$, $H_{\gamma_{i}}=\otimes_{j=1}^{k_{i}}H_{n_{ij}}$, $\gamma_{i}=\{n_{i1},n_{i2},\cdots,n_{ik_{i}}\}$, $n_{i1}<n_{i2}<\cdots<n_{ik_{i}}$; otherwise, $|\psi\rangle\langle\psi|$ is called $k$-nonseparable. A mixed state $\rho\in\mathcal{S}(H)$ is $k$-separable if $\rho$ can be written as $\rho=\sum_{r}p_{r}|\psi_{r}\rangle\langle\psi_{r}|$, where each $|\psi_{r}\rangle\langle\psi_{r}|\in{\mathcal{S}}(H)$ is a $k$-separable pure state and $p_{r}\geq 0$ with $\sum_{r}p_{r}=1$; otherwise, $\rho$ is $k$-nonseparable. Particularly, $\rho$ is fully separable if $\rho$ is $N$-separable; and is genuine entangled if $\rho$ is 2-nonseparable.
On the other hand, a pure state $|\psi\rangle\langle\psi|\in\mathcal{S}(H)$ is $k$-producible if it can be represented as
$|\psi\rangle=\otimes_{i=1}^{m}|\psi_{\gamma_{i}}\rangle,$
where $\gamma_{1}|\gamma_{2}|\cdots|\gamma_{m}$ is a $m$-partition of $\{1,2,\cdots,N\}$, and the number of particles in each subset $\gamma_{i}$ is not greater than $k$. For any mixed state $\rho$, we say that $\rho$ is $k$-producible if it can be represented as a convex combination of $k$-producible pure states; otherwise, we say that $\rho$ contains $(k+1)$-partite entanglement GOTG .
Denote by $\mathcal{S}_{k}$ the set of all $k$-separable quantum states ($2\leq k\leq N$) and by $\mathcal{P}_{k}$ the set of all $k$-producible quantum states ($1\leq k\leq N-1$). By the above definitions, it is easily seen that, if a state is $k$-separable, then it is also (k1)(k-1)-separable; and if a state is $k$-producible, then it is also $(k+1)$-producible. That is,
$\mathcal{S}_{N}\subset\mathcal{S}_{N-1}\subset\cdots\subset\mathcal{S}_{2}\subset\mathcal{S}_{1};\ \ \mathcal{P}_{N}\supset\mathcal{P}_{N-1}\supset\cdots\mathcal{P}_{2}\supset\mathcal{P}_{1},$
where both $\mathcal{S}_{N}$ and $\mathcal{P}_{1}$ stand for the set of all fully separable states, while $\mathcal{S}_{1}=\mathcal{P}_{N}=\mathcal{S}(H)$.
II.2 Wigner-Yanase skew information
In the study of measurement theory, Wigner and Yanase introduced the Wigner-Yanase skew information from an information-theoretic perspective. For a quantum state $\rho\in{\mathcal{S}}(H)$, the Wigner-Yanase skew information of $\rho$ with respect to an observable $A$ is defined by
$I(\rho,A)=-\frac{1}{2}(\text{tr}[\rho^{1/2},A]^{2})$
with [A,B]=ABBA[A,B]=AB-BA as usual WY . This quantity $I(\rho,A)$ represents the amount of information about the values of observables that do not commute with $A$, which is skew-symmetric with respect to $A$.
II.3 Quantum Fisher information
The quantum Fisher information of a quantum state $\rho\in{\mathcal{S}}(H)$ with respect to an observable $A\in{\mathcal{B}}(H)$ is defined as HWC
$F(\rho,A)=\frac{1}{4}{\rm tr}(\rho L^{2}),$
where $L$ is the symmetric logarithmic derivative operator defined via
$i(\rho A-A\rho)=\frac{1}{2}(\rho L+L\rho).$
If $\rho$ has the spectral decomposition $\rho=\sum_{l}\lambda_{l}|l\rangle\langle l|$, then $F(\rho,A)$ can be expressed as
$F(\rho,A)=\sum_{l,l^{\prime}}\frac{\left(\lambda_{l}-\lambda_{l^{\prime}}\right)^{2}}{2\left(\lambda_{l}+\lambda_{l^{\prime}}\right)}\left|\left<l|A|l^{\prime}\right>\right|^{2}.$
II.4 Generalized Wigner-Yanase skew information
For any $s\in(-\infty,0)$, define a binary function $f_{s}$ on $\mathbb{R}_{+}\times\mathbb{R}_{+}$ ($\mathbb{R}_{+}=[0,+\infty)$) as follows:
$f_{s}(a,b)=\begin{cases}\left(\frac{a^{s}+b^{s}}{2}\right)^{1/s}&\text{if}\ a>0,b>0,\\ 0&\text{if}\ a=0\ \text{or}\ b=0,\end{cases}$
which represents the generalized mean of two positive real numbers $a$ and $b$ with equal weights. In addition, define
$f_{0}(a,b):=\lim_{s\to 0}f_{s}(a,b)=\sqrt{ab}$
and
$f_{-\infty}(a,b):=\lim_{s\to-\infty}f_{s}(a,b)=\min\{a,b\}.$
For any quantum state $\rho\in{\mathcal{S}}(H)$ with the spectral decomposition $\rho=\sum_{l}\lambda_{l}|\psi_{l}\rangle\langle\psi_{l}|$, the generalized Wigner-Yanase skew information about an observable $A$ is given by YQ
$I^{s}(\rho,A)$= $\text{tr}\left(\rho A^{2}\right)-\sum_{l,l^{\prime}}f_{s}\left(\lambda_{l},\lambda_{l^{\prime}}\right)\left|\left<\psi_{l}|A|\psi_{l^{\prime}}\right>\right|^{2}$
$=$ $\sum_{l\neq l^{\prime}}\left[\lambda_{l}-f_{s}\left(\lambda_{l},\lambda_{l^{\prime}}\right)\right]\left|\left<\psi_{l}|A|\psi_{l^{\prime}}\right>\right|^{2}.$
Particularly, when s=1s=-1 and $s=0$, $I^{-1}(\rho,A)$ and $I^{0}(\rho,A)$ reduce to the quantum Fisher information and the Wigner-Yanase skew information, respectively.
The authors in YQ proved that the generalized Wigner-Yanase skew information has the following properties:
(1) Monotonicity: $I^{s}(\rho,A)$ decreases monotonically with respect to $s$, that is,
$I^{0}(\rho,A)\leq I^{-1}(\rho,A)\leq\cdots\leq I^{-\infty}(\rho,A)\leq V(\rho,A),$
where $V(\rho,A)=\text{Tr}(\rho A^{2})-[\text{Tr}(\rho A)]^{2}$ stands for the variance. Moreover, the equalities hold if $\rho$ is pure.
(2) Convexity: For any $s$, $I^{s}$ is convex, that is,
$I^{s}\left(\sum_{i}p_{i}\rho_{i},A\right)\leq\sum_{i}p_{i}I^{s}\left(\rho_{i},A\right),$
where $\rho_{i}\in{\mathcal{S}}(H)$, $p_{i}\geq 0$ and $\sum_{i}p_{i}=1$.
(3) Additivity: For any $\rho_{i}\in{\mathcal{S}}(H_{i})$ and any observable $A_{i}\in\mathcal{B}(H_{i})$, $i=1,2,\cdots,N$, we have
$I^{s}\left(\otimes_{i=1}^{N}\rho_{i},\sum_{i=1}^{N}{\bf A_{i}}\right)=\sum_{i=1}^{N}I^{s}\left(\rho_{i},A_{i}\right),$
where ${\bf A_{i}}:=(\otimes_{j=1}^{i-1}\mathbb{I}_{j})\otimes A_{i}\otimes(\otimes_{j=i+1}^{N}\mathbb{I}_{j})$.
II.5 Mutually unbiased measurements (MUMs)
Let $P^{(b)}=\{P_{n}^{(b)}:n=1,\cdots,d\}$ be two positive-operator-valued measurements (POVMs), $b=1,2$. $P^{(1)}$ and $P^{(2)}$ are called MUMs if they satisfy the following conditions:
$\begin{cases}\mathrm{tr}(P_{n}^{(b)})=1,\\ \mathrm{tr}(P_{n}^{(b)}P_{n^{\prime}}^{(b^{\prime})})=&\delta_{n,n^{\prime}}\delta_{b,b^{\prime}}\kappa\\ &+(1-\delta_{n,n^{\prime}})\delta_{b,b^{\prime}}\frac{1-\kappa}{d-1}+(1-\delta_{b,b^{\prime}})\frac{1}{d},\end{cases}$
where $b,b^{\prime}\in\{1,2\},\ n,n^{\prime}\in\{1,\cdots,d\}$, and the parameter $\kappa$ satisfies $\frac{1}{d}<\kappa\leq 1$. When $\kappa=1$, it reduces to mutually unbiased bases (MUBs) WF .
Kalev and Gour in KG proposed a construction method for MUMs in any dimensional quantum system. Take an orthonormal basis $\{F_{k}\}_{k=1}^{d^{2}-1}$ consisting of Hermitian and traceless operators acting on $H$. Arrange these basis elements on a grid of d1d-1 columns and $d+1$ rows, and denote them as $\{F_{n,b}\}$ with $n=1,2,\ldots,d-1$ and $b=1,2,\ldots,d+1$. Define $d(d+1)$ operators $\{F_{n}^{(b)}\}$ by
$F_{n}^{(b)}=\left\{\begin{array}[]{ll}F^{(b)}-(d+\sqrt{d})F_{n,b}&\text{for}\ n=1,\cdots,d-1,\\ (1+\sqrt{d})F^{(b)}&\text{for}\ n=d,\end{array}\right$. (1)
where $F^{(b)}=\sum_{n=1}^{d-1}F_{n,b}$, $b=1,2,\cdots,d+1$. It is a direct calculation that
$\sum_{n=1}^{d}F_{n}^{(b)}=0,\ \ \forall b\in\{1,2,\cdots,d+1\}.$ (2)
The measurement operators $\{P_{n}^{(b)}\}$ are given by:
$P_{n}^{(b)}=\frac{1}{d}\mathbb{I}+tF_{n}^{(b)},$ (3)
where the free parameter $t$ is chosen such that $P_{n}^{(b)}\geq 0$; and
$\kappa=\frac{1}{d}+t^{2}(1+\sqrt{d})^{2}(d-1).$ (4)
Then $\{P^{(b)}=\{P_{n}^{(b)}\}\}$ form a complete set of $d+1$ mutually unbiased measurements, where $b$ labels the measurement and $n$ labels the outcome. Moreover, any complete set of MUMs has this form.
It was also proven in LL that $\left\{\frac{\mathbb{I}}{\sqrt{d}},F_{n,b}:n=1,\cdots,d-1;b=1,\cdots,d+1\right\}$ forms a set of local orthogonal observables (LOOs) on $H$; and any complete set of MUMs $\{P_{n}^{(b)}:n=1,2,\cdots,d;b=1,2,\cdots,d+1\}$ must satisfy the following equation:
$\sum_{b=1}^{d+1}\sum_{n=1}^{d}(P_{n}^{(b)})^{2}=(d+1)\kappa\mathbb{I}.$ (5)
III $k$-separability and $k$-partite entanglement detection criteria
In this section, we will establish some entanglement detection criteria for $k$-separability and $k$-partite entanglement in multipartite quantum systems.
For convenience, assume $H_{i}=H_{0}$ with $\dim H_{0}=d$, $i=1,2,\cdots,N$. Take the same set of $d+1$ MUMs acting on each subsystem $H_{i}$
$\{P^{(i,b)}=\{P_{i,n}^{(b)}\}:b=1,2,\cdots,d+1;n=1,2,\cdots,d\},$
where $i$ is used to label the subsystem on which the operators act, $i=1,2,\cdots,N$. For any subset $\Gamma\subset\{1,2,\cdots,N\}$ with $\sharp(\Gamma)$ the number of the elements in $\Gamma$, write
${\bf P}_{\Gamma,n}^{(b)}=\sum_{i\in\Gamma}{\bf P}_{i,n}^{(b)}$
with
${\bf P}_{i,n}^{(b)}=(\otimes_{j=1}^{i-1}\mathbb{I}_{j})\otimes P_{i,n}^{(b)}\otimes(\otimes_{j=i+1}^{\sharp(\Gamma)}\mathbb{I}_{j});$
and in particular, when $\Gamma=\{1,2,\cdots,N\}$, write
${\bf P}_{N,n}^{(b)}=\sum_{i=1}^{N}{\bf P}_{i,n}^{(b)}.$
For ${\bf P}_{\Gamma,n}^{(b)}$, we have the following result.
Proposition 1.
For any $p\neq q\in\Gamma$, we have
$\sum_{b=1}^{d+1}\sum_{n=1}^{d}{\bf P}_{p,n}^{(b)}{\bf P}_{q,n}^{(b)}\leq(1+\kappa)(\otimes_{i\in\Gamma}\mathbb{I}_{i}).$
Proof. Without loss of generality, assume $p<q$. By the definition, one has
${\bf P}_{p,n}^{(b)}{\bf P}_{q,n}^{(b)}=(\otimes_{j=1}^{p-1}\mathbb{I}_{j})\otimes P_{p,n}^{(b)}\otimes(\otimes_{j=p+1}^{q-1}\mathbb{I}_{j})\otimes P_{q,n}^{(b)}\otimes(\otimes_{j=q+1}^{\sharp(\Gamma)}\mathbb{I}_{j}).$
It follows from Eqs.(1)-(5) that
$\sum_{b=1}^{d+1}\sum_{n=1}^{d}{\bf P}_{p,n}^{(b)}{\bf P}_{q,n}^{(b)}$
$=$ $\sum_{b=1}^{d+1}\sum_{n=1}^{d}(\otimes_{j=1}^{p-1}\mathbb{I}_{j})\otimes(\frac{1}{d}\mathbb{I}_{p}+tF_{n}^{(b)})\otimes(\otimes_{j=p+1}^{q-1}\mathbb{I}_{j})\otimes(\frac{1}{d}\mathbb{I}_{q}+tF_{n}^{(b)})\otimes(\otimes_{j=q+1}^{\sharp(\Gamma)}\mathbb{I}_{j})$
$=$ $\sum_{b=1}^{d+1}\sum_{n=1}^{d}\big{[}\frac{(\otimes_{i\in\Gamma}\mathbb{I}_{i})}{d^{2}}+\frac{t}{d}((\otimes_{j=1}^{p-1}\mathbb{I}_{j})\otimes F_{n}^{(b)}\otimes(\otimes_{j=p+1}^{\sharp(\Gamma)}\mathbb{I}_{j})+(\otimes_{j=1}^{q-1}\mathbb{I}_{j})\otimes F_{n}^{(b)}\otimes(\otimes_{j=q+1}^{\sharp(\Gamma)}\mathbb{I}_{j}))$
$+t^{2}(\otimes_{j=1}^{p-1}\mathbb{I}_{j})\otimes F_{n}^{(b)}\otimes(\otimes_{j=p+1}^{q-1}\mathbb{I}_{j})\otimes F_{n}^{(b)}\otimes(\otimes_{j=q+1}^{\sharp(\Gamma)}\mathbb{I}_{j})\big{]}$
$=$ $\frac{d(d+1)}{d^{2}}(\otimes_{i\in\Gamma}\mathbb{I}_{i})+t^{2}\sum_{b=1}^{d+1}\sum_{n=1}^{d}(\otimes_{j=1}^{p-1}\mathbb{I}_{j})\otimes F_{n}^{(b)}\otimes(\otimes_{j=p+1}^{q-1}\mathbb{I}_{j})\otimes F_{n}^{(b)}\otimes(\otimes_{j=q+1}^{\sharp(\Gamma)}\mathbb{I}_{j})$
$=$ $\frac{d+1}{d}(\otimes_{i\in\Gamma}\mathbb{I}_{i})+t^{2}(d+\sqrt{d})^{2}\sum_{b=1}^{d+1}\sum_{n=1}^{d-1}(\otimes_{j=1}^{p-1}\mathbb{I}_{j})\otimes F_{n,b}\otimes(\otimes_{j=p+1}^{q-1}\mathbb{I}_{j})\otimes F_{n,b}\otimes(\otimes_{j=q+1}^{\sharp(\Gamma)}\mathbb{I}_{j})$
$\leq$ $\frac{d+1}{d}(\otimes_{i\in\Gamma}\mathbb{I}_{i})+t^{2}(d+\sqrt{d})^{2}(1-\frac{1}{d})(\otimes_{i\in\Gamma}\mathbb{I}_{i})=(1+\kappa)(\otimes_{i\in\Gamma}\mathbb{I}_{i}).$
$\Box$
Proposition 2.
Assume that $\Gamma\subset\{1,2,\cdots,N\}$ is any subset. For any $\rho\in{\mathcal{S}}(H_{1}\otimes H_{2}\otimes\cdots\otimes H_{N})$, denoting by $\rho_{\Gamma}$ the reduced state of $\rho$ on the subsystem $\otimes_{i\in\Gamma}H_{i}$, we have
$\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho_{\Gamma},{\bf P}_{\Gamma,n}^{(b)})\leq\sharp(\Gamma)^{2}(\kappa-\frac{1}{d})+\sharp(\Gamma)(d\kappa-1).$
Proof. For any subset $\Gamma\subset\{1,2,\cdots,N\}$, by Proposition 3.1 and Eq.(5), we have
$\sum_{b=1}^{d+1}\sum_{n=1}^{d}\big{(}{\bf P}_{\Gamma,n}^{(b)}\big{)}^{2}$
$=\sum_{b=1}^{d+1}\sum_{n=1}^{d}\big{(}\sum_{i\in\Gamma}{\bf P}_{i,n}^{(b)}\big{)}^{2}$
$=\sum_{b=1}^{d+1}\sum_{n=1}^{d}\big{(}\sum_{i\in\Gamma}({\bf P}_{i,n}^{(b)})^{2}+\sum_{p\not=q\in\Gamma}{\bf P}_{p,n}^{(b)}{\bf P}_{q,n}^{(b)}\big{)}$
$=\sum_{i\in\Gamma}\sum_{b=1}^{d+1}\sum_{n=1}^{d}((\otimes_{j=1}^{i-1}\mathbb{I}_{j})\otimes P_{i,n}^{(b)}\otimes(\otimes_{j=i+1}^{\sharp(\Gamma)}\mathbb{I}_{j}))^{2}$
$+\sum_{p\neq q\in\Gamma}\sum_{b=1}^{d+1}\sum_{n=1}^{d}{\bf P}_{p,n}^{(b)}{\bf P}_{q,n}^{(b)}$
$\leq\sharp(\Gamma)(d+1)\kappa(\otimes_{i\in\Gamma}\mathbb{I}_{i})+(\sharp(\Gamma)^{2}-\sharp(\Gamma))(1+\kappa)(\otimes_{i\in\Gamma}\mathbb{I}_{i})$
$=[\sharp(\Gamma)^{2}(1+\kappa)+\sharp(\Gamma)(d\kappa-1)](\otimes_{i\in\Gamma}\mathbb{I}_{i}).$
On the other hand, by Eq.(2), one has
$\sum_{b=1}^{d+1}\sum_{n=1}^{d}\big{(}\mathrm{tr}\rho_{\Gamma}{\bf P}_{\Gamma,n}^{(b)}\big{)}^{2}$
$=\sum_{b=1}^{d+1}\sum_{n=1}^{d}[\mathrm{tr}\rho_{\Gamma}\sum_{i\in\Gamma}{\bf P}_{i,n}^{(b)}]^{2}=\sum_{b=1}^{d+1}\sum_{n=1}^{d}[\sum_{i\in\Gamma}\mathrm{tr}\rho_{\Gamma}{\bf P}_{i,n}^{(b)}]^{2}$
$=\sum_{b=1}^{d+1}\sum_{n=1}^{d}[\sum_{i\in\Gamma}\mathrm{tr}\rho_{\Gamma}(\frac{\otimes_{i\in\Gamma}\mathbb{I}_{i}}{d}+t{\bf F}_{i,n}^{(b)})]^{2}$
$=\sum_{b=1}^{d+1}\sum_{n=1}^{d}(\frac{\sharp(\Gamma)}{d}+t\sum_{i\in\Gamma}\mathrm{tr}\rho_{\Gamma}{\bf F}_{i,n}^{(b)})^{2}$
$=\sum_{b=1}^{d+1}\sum_{n=1}^{d}\left(\frac{\sharp(\Gamma)}{d}\right)^{2}+t^{2}\sum_{b=1}^{d+1}\sum_{n=1}^{d}\left(\sum_{i\in\Gamma}\mathrm{tr}\left(\rho_{\Gamma}{\bf F}_{i,n}^{(b)}\right)\right)^{2}$
$\geq\sum_{b=1}^{d+1}\sum_{n=1}^{d}(\frac{\sharp(\Gamma)}{d})^{2}=\sharp(\Gamma)^{2}(1+\frac{1}{d}),$
where ${\bf F}_{i,n}^{(b)}=(\otimes_{j=1}^{i-1}\mathbb{I}_{j})\otimes F_{i,n}^{(b)}\otimes(\otimes_{j=i+1}^{\sharp(\Gamma)}\mathbb{I}_{j})$. Combining the above two inequalities gives
$\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho_{\Gamma},{\bf P}_{\Gamma,n}^{(b)})$
$\leq\sum_{b=1}^{d+1}\sum_{n=1}^{d}\Big{[}\mathrm{tr}\,\rho_{\Gamma}({\bf P}_{\Gamma,n}^{(b)})^{2}-(\mathrm{tr}\,\rho_{\Gamma}{\bf P}_{\Gamma,n}^{(b)})^{2}\Big{]}$
$\leq[\sharp(\Gamma)^{2}(1+\kappa)+\sharp({\Gamma})(d\kappa-1)]-\sharp(\Gamma)^{2}(1+\frac{1}{d})$
$=\sharp(\Gamma)^{2}(\kappa-\frac{1}{d})+\sharp(\Gamma)(d\kappa-1).$
$\Box$
Now, we can give a $k$-nonseparability criterion and a $k$-partite entanglement criterion by generalized Wigner-Yanase skew informations and MUMs.
Theorem 3.
If $\rho$ is $k$-separable, then the following inequality holds:
$\sup_{s\in[-\infty,0]}\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})$
$=$ $\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{-\infty}\left(\rho,\mathbf{P}_{N,n}^{(b)}\right)$
$\leq$ $N(d\kappa-1)+\left[(N-k+1)^{2}+k-1\right](\kappa-\frac{1}{d}).$
Proof. Define the generalized skew information function:
$g_{s}(\rho)=\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}\left(\rho,\mathbf{P}_{N,n}^{(b)}\right),\quad s\in[-\infty,0].$
Since $\rho$ is $k$-separable, then $\rho$ can be written as a convex combination of $k$-separable pure states, that is, $\rho=\sum_{r}p_{r}\rho_{r}$, where $\rho^{(r)}$ has the form
$\rho_{r}=\rho^{(r)}_{\Gamma_{1}}\otimes\rho^{(r)}_{\Gamma_{2}}\otimes\cdots\otimes\rho^{(r)}_{\Gamma_{k}}$
for some $k$-partition $\Gamma_{1}|\Gamma_{2}|\ldots|\Gamma_{k}$ of $\{1,2,\cdots,N\}$. By Proposition 2, and by the convexity and additivity of the generalized Wigner-Yanase skew information, we have
$g_{s}(\rho)=$ $\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})$
$\leq$ $\sum_{b=1}^{d+1}\sum_{n=1}^{d}\sum_{r}p_{r}I^{s}(\rho_{r},{\bf P}_{N,n}^{(b)})$
$=$ $\sum_{b=1}^{d+1}\sum_{n=1}^{d}\sum_{r}p_{r}\sum_{t=1}^{k}I^{s}(\rho_{\Gamma_{t}}^{(r)},{\bf P}_{\Gamma_{t},n}^{(b)})$
$\leq$ $\sum_{r}p_{r}\sum_{i=1}^{k}[\sharp(\Gamma_{i})^{2}(\kappa-\frac{1}{d})+\sharp(\Gamma_{i})(d\kappa-1)]$
$=$ $(\kappa-\frac{1}{d})\sum_{i=1}^{k}\sharp(\Gamma_{i})^{2}+N(d\kappa-1).$
The next goal is to maximize $\sum_{i=1}^{k}\sharp(\Gamma_{i})^{2}$. To do this, write
$x=(\sharp(\Gamma_{i_{1}}),\sharp(\Gamma_{i_{2}}),\dots,\sharp(\Gamma_{i_{k}}))^{t}\in\mathbb{R}^{k},$
where $\sharp(\Gamma_{i_{j}})\in\mathbb{Z}^{+}$ (the set of positive integers), $\sharp(\Gamma_{i_{1}})\geq\sharp(\Gamma_{i_{2}})\geq\dots\geq\sharp(\Gamma_{i_{k}})$ with $\sum_{j=1}^{k}\sharp(\Gamma_{i_{j}})=N$, $i_{j}\not=i_{l}\in\{1,2,\cdots,k\}$. Denote by $X\subset\mathbb{R}^{k}$ the set consisting of all the above vectors $x$. Write
$y_{0}=(N-k+1,1,\cdots,1)^{t}\in\mathbb{R}^{k}.$
It is obvious that $y_{0}\in X$. Recall that, for any vectors $x=(x_{1},x_{2},\cdots,x_{k})^{t}$ and $y=(y_{1},y_{2},\cdots,y_{k})^{k}\in{\mathbb{R}}^{k}$, $x\prec y$ means that $\sum_{i=1}^{m}x_{i}\leq\sum_{i=1}^{m}y_{i}$ holds for all $m=1,2,\cdots,k-1$ and $\sum_{i=1}^{k}x_{i}=\sum_{i=1}^{k}y_{i}$. Then we have $x\prec y_{0}$ for all $x\in X$; otherwise, if there exists some $m^{\prime}$ such that $\sum_{i=1}^{m^{\prime}}x_{i}>\sum_{i=1}^{m^{\prime}}y_{i}=N-k+m^{\prime},$ then there must be some $x_{0}=0$, a contradiction. Note that the function $x\mapsto\sum_{i=1}^{k}x_{i}^{2}$ is Schur convex. It follows that $\sum_{i=1}^{k}x_{i}^{2}\leq(N-k+1)^{2}+(k-1)$. This implies that $\sum_{i=1}^{k}\sharp(\Gamma_{i})^{2}$ attains the maximum at the partition $\{\Gamma_{1},\cdots,\Gamma_{k}\}$ with $\sharp(\Gamma_{1})=N-k+1,\ \sharp(\Gamma_{i})=1$ for all $i=2,3,\cdots,k$.
Hence, for any $s\in[-\infty,0]$, one obtains
$g_{s}(\rho)\leq N(d\kappa-1)+\left[(N-k+1)^{2}+k-1\right]\left(\kappa-\frac{1}{d}\right).$
Now, by the monotonicity of the generalized Wigner-Yanase skew information, we achieve
$\begin{array}[]{rl}&\sup_{s\in[-\infty,0]}g_{s}(\rho)=g_{-\infty}(\rho)\\ =&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{-\infty}\left(\rho,\mathbf{P}_{N,n}^{(b)}\right)\\ \leq&N(d\kappa-1)+\left[(N-k+1)^{2}+k-1\right]\left(\kappa-\frac{1}{d}\right),\end{array}$
as desired. $\Box$
Theorem 4.
If $\rho$ is $k$-producible, then the following inequality holds:
$\begin{array}[]{rl}&\sup_{s\in[-\infty,0]}\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})=\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{-\infty}(\rho,{\bf P}_{N,n}^{(b)})\\ \leq&\begin{cases}Nk(\kappa-\frac{1}{d})+N(d\kappa-1)&{\rm if}\ \ N=pk;\\ (N^{2}+p^{2}k^{2}+pk^{2}-2pkN)(\kappa-\frac{1}{d})+N(d\kappa-1)&{\rm if}\ \ 0<N-pk<k,\end{cases}\end{array}$
where $p=[\frac{N}{k}]$.
Proof. Since $\rho$ is $k$-producible, $\rho$ can be written as $\rho=\sum_{r}p_{r}\rho_{r}$, where
$\rho_{r}=\rho^{(r)}_{\Gamma_{1}}\otimes\rho^{(r)}_{\Gamma_{2}}\otimes\cdots\otimes\rho^{(r)}_{\Gamma_{l_{r}}}$
for some $l_{r}$-partition $\Gamma_{1}|\Gamma_{2}|\ldots|\Gamma_{l_{r}}$ with $\sharp(\Gamma_{i})\leq k$ for every $i=1,\cdots,l_{r}\leq N$. Let $\sharp(\Gamma_{i})=0$ for $i=l_{r}+1,\ldots,N$. Then, by Proposition 3.2, we have
$\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})$
$\leq\sum_{b=1}^{d+1}\sum_{n=1}^{d}\sum_{r}p_{r}I^{s}(\rho_{r},{\bf P}_{N,n}^{(b)})$
$=\sum_{b=1}^{d+1}\sum_{n=1}^{d}\sum_{r}p_{r}\sum_{t=1}^{l_{r}}I^{s}(\rho_{\Gamma_{t}}^{(r)},{\bf P}_{\Gamma_{t},n}^{(b)})$
$\leq\sum_{r}p_{r}\sum_{i=1}^{l_{r}}[\sharp(\Gamma_{i})^{2}(\kappa-\frac{1}{d})+\sharp(\Gamma_{i})(d\kappa-1)]$
$=(\kappa-\frac{1}{d})\sum_{i=1}^{N}\sharp(\Gamma_{i})^{2}+N(d\kappa-1).$
Write
$x=(\sharp(\Gamma_{i_{1}}),\sharp(\Gamma_{i_{2}}),\dots,\sharp(\Gamma_{i_{l_{r}}}),0,\ldots,0)^{t}\in\mathbb{R}^{N},$
where $\sharp(\Gamma_{i_{j}})\in\mathbb{Z}^{+}$ with $\sharp(\Gamma_{i_{j}})\leq k$, $\sharp(\Gamma_{i_{1}})\geq\sharp(\Gamma_{i_{2}})\geq\dots\geq\sharp(\Gamma_{i_{l_{r}}})$ with $\sum_{j=1}^{l_{r}}\sharp(\Gamma_{i_{j}})=N$, $i_{j}\not=i_{p}\in\{1,2,\cdots,l_{r}\}$. Denote by $Y\subset\mathbb{R}^{N}$ the set consisting of all the above vectors $x$. Write
$y_{0}=(\underbrace{k,k,\dots,k}_{p},N-pk,0,\dots,0)^{t}\in\mathbb{R}^{N},$
where $p=[\frac{N}{k}]$ the integer part of $\frac{N}{k}$. It is easily seen that $y_{0}\in Y$ and $x\prec y_{0}$ for all $x\in Y$. By the Schur convexity of the function $x\mapsto\sum_{i=1}^{N}x_{i}^{2}$ again, we see that $\sum_{i=1}^{N}t_{i}^{2}$ attains the maximum at $y_{0}$. Consequently, $\sum_{i=1}^{N}\sharp{(\Gamma_{i})}^{2}$ achieves the maximum at the $(p+1)$-partition $\{\Gamma_{1},\cdots,\Gamma_{l}\}$, where $l=p$ if $N=pk$ and $l=p+1$ if $N\not=pk$, with
$\sharp(\Gamma_{1})=\sharp(\Gamma_{2})=\dots=\sharp(\Gamma_{p})=k,\quad\sharp(\Gamma_{p+1})=N-pk.$
Thus, if $N=pk$, then
$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ &\leq pk^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)=Nk(\kappa-\frac{1}{d})+N(d\kappa-1);\end{array}$
and if $N=pk+m$ with $0<m<k$, then
$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ &\leq pk^{2}(\kappa-\frac{1}{d})+(N-pk)^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)\\ &=(N^{2}+p^{2}k^{2}+pk^{2}-2pkN)(\kappa-\frac{1}{d})+N(d\kappa-1).\end{array}$
Thus, the monotonicity of generalized Wigner-Yanase skew information ensures that the tightest bounds are achieved in the limiting case $s=$-$\infty$. This completes the proof of the theorem. $\Box$
Consequently, any violation of the inequalities in Theorems 3-4 indicate $k$-nonseparability and $(k+1)$-partite entanglement, respectively.
To verify the validity of these criteria, we provide several examples below.
Example 5.
Consider the 6-partite quantum state
$\rho(p)=p|\psi\rangle\langle\psi|+\frac{1-p}{64}\mathbb{I},$
where $|\psi\rangle=\frac{1}{\sqrt{6}}(|000001\rangle+|000010\rangle+|000100\rangle+|001000\rangle+|010000\rangle+|100000\rangle)$.
In this case, we illustrate the situation by taking s=1,0,s=-1,0,-$\infty$, respectively. According to Theorem 3 and Theorem 4, we can obtain the following conclusions: when $p_{k}<p\leq 1$, $\rho(p)$ is $k$-nonseparable; when $\tilde{p_{k}}<p\leq 1$, $\rho(p)$ has $(k+1)$-partite entanglement. In HQGY , the authors showed that $\rho(p)$ is $k$-nonseparable if $p_{k}^{\prime}<p\leq 1$, and $\rho(p)$ contains $(k+1)$-partite entanglement if $\tilde{p_{k}^{\prime}}<p\leq 1$. See the following tables.
$k$ 2 3 4 5 6
$p_{k}$ 0.6523 0.5211 0.4225 0.3567 0.3237
$p_{k}^{\prime}$ $\backslash$ 0.5816 0.4433 0.3443 0.2649
Table 1: Range for detecting $k$-nonseparability for s=1s=-1.
$k$ 1 2 3 4 5
$\tilde{p_{k}}$ 0.3237 0.4225 0.5211 0.5539 0.6523
$\tilde{p_{k}^{\prime}}$ 0.2649 0.5026 0.6210 0.6605 0.7591
Table 2: Range for detecting $(k+1)$-partite entanglement for s=1s=-1.
$k$ 2 3 4 5 6
$p_{k}$ 0.9943 0.9927 0.9913 0.9903 0.9898
$p_{k}^{\prime}$ $\backslash$ 0.9990 0.9978 0.9961 0.9939
Table 3: Range for identifying $k$-nonseparability for $s=0$.
$k$ 1 2 3 4 5
$\tilde{p_{k}}$ 0.9898 0.9913 0.9927 0.9931 0.9943
[…]
Table 4: Range for identifying $(k+1)$-partite entanglement for $s=0$.
$k$ 2 3 4 5 6
$p_{k}$ 0.3958 0.3125 0.25 0.2083 0.1875
$p_{k}^{\prime}$ $\backslash$ 0.75 0.625 0.5 0.375
Table 5: Range for identifying $k$-nonseparability for $s=$-$\infty$.
$k$ 1 2 3 4 5
$\tilde{p_{k}}$ 0.1875 0.25 0.3125 0.3333 0.3958
[…]
Table 6: Range for identifying $(k+1)$-partite entanglement for $s=$-$\infty$.
As clearly demonstrated in Tables I-VI, the case $s=$-$\infty$ exhibits the widest detection range, consistently outperforming both s=1s=-1 and $s=0$ in identifying entangled states across all tested scenarios. This observation further corroborates our theoretical finding that the optimal detection capability is achieved precisely at the limiting case $s=$-$\infty$, where the generalized Wigner-Yanase skew information attains its maximal sensitivity. The enhanced performance at $s=$-$\infty$ can be attributed to the complete utilization of all high-order quantum coherences in the state detection.
In addition, it can be seen that, Theorem 3 can detect more $k$-nonseparable quantum states; and, except for $k=1$, Theorem 4 can detect more $(k+1)$-partite entangled quantum states, which indicate that our criteria have better effectiveness. $\Box$
Example 6.
Consider an 11-partite quantum state
$\rho=p|\psi\rangle\langle\psi|+\frac{1-p}{2^{11}}\mathbb{I}$
where $|\psi\rangle=\frac{1}{\sqrt{2}}(|0\rangle^{\otimes 11}+|1\rangle^{\otimes 11})$.
In this example, we take $s=$-$\infty$ for an illustration. According to Theorem 3 and Theorem 4, we can conclude that $\rho(p)$ is $k$-nonseparable if $p_{k}<p\leq 1$ and exhibits $(k+1)$-partite entanglement if $\tilde{p_{k}}<p\leq 1$. Note that, $p_{k}^{\prime}$ is the result obtained using the method in HLS , and $\tilde{p_{k}^{\prime}}$ is obtained in HQGY , as shown in the Tables VII-VIII.
$k$ 2 3 4 5 6 7 8 9 10 11
$p_{k}$ 0.4300 0.3671 0.3111 0.2622 0.2202 0.1853 0.1573 0.1363 0.1223 0.1153
$p_{k}^{\prime}$ 0.8532 0.7205 0.6017 0.4969 0.4061 0.3293 0.2664 0.2175 0.1826 0.1546
Table 7: Range for identifying $k$-nonseparability.
$k$ 1 2 3 4 5 6 7 8 9 10
$\tilde{p_{k}}$ 0.1153 0.1503 0.1853 0.2202 0.2552 0.2902 0.3041 0.3321 0.3741 0.4300
$\tilde{p_{k}^{\prime}}$ 0.0009 0.0312 0.1248 0.2498 0.2498 0.4997 0.4997 0.4997 0.4997 0.4997
Table 8: Range for identifying $(k+1)$-partite entanglement.
From the tables, it can be seen that, for $2\leq k\leq 11$, Theorem 3 can identify more $k$-nonseparable quantum states for all $k$ compared to that in HLS , which reveals that our criterion has better effectiveness. For $k=4,6,7,8,9,10$, utilizing Theorem 4 can detect more $(k+1)$-partite entangled quantum states compared to that in HQGY2 , which demonstrates that our criterion performs better for biger $k$s. $\Box$
Finally, we give an example of applying Theorem 4 to distinct the networks by detecting the depth of quantum networks.
A quantum network is a system that connects quantum devices to share quantum information, enabling tasks like secure communication and distributed computing using quantum properties LMX ; KHJ . Generally, a quantum network consists of $n$ nodes (subsystems, denote by $A_{j}$, $j=1,2,\cdots,n$) and $r$ quantum resources $\rho_{i}$ shared by some of the nodes. Let $L_{i}\subseteq\{1,2,\cdots,n\}$ the subset of index of the nodes that $\{A_{j}:j\in L_{i}\}$ is exact the subset of all nodes that share the $i$-th resource $\rho_{i}$. Denote by $\rho_{i}=\rho_{L_{i}}$, $i=1,2,\cdots,r$. Then the corresponding quantum network state has the form:
$\rho=\rho_{L_{1}}\otimes\rho_{L_{2}}\otimes\cdots\rho_{L_{r}}.$
Accordingly, for every $j=1,2,\ldots,n$, the node $A_{j}$ may share several $\rho_{L_{i}}$ with other nodes, that is $j\in L_{i}$. Let $F_{j}=\{i_{j}:A_{j}\in L_{i}\}\subseteq\{1,2,\ldots,r\}$. Then the subsystem hold by $A_{j}$ is a $\sharp F_{j}$-partite composite system $H_{A_{j}}=\bigotimes_{i_{j}\in F_{j}}H_{A_{j_{i}}}$. Every quantum network state determined by this network has the form:
$\rho=\sum_{\lambda\in\Lambda}p_{\Lambda}(\lambda)(\Omega_{1}^{\lambda}\otimes\cdots\otimes\Omega_{n}^{\lambda})(\rho_{L_{1}}^{\lambda}\otimes\rho_{L_{2}}^{\lambda}\otimes\cdots\rho_{L_{r}}^{\lambda}),$ (6)
where $\{p_{\Lambda}(\lambda)\}$ is a probability distribution, $\Omega_{j}^{\lambda}$ is a quantum channel of the subsystem $A_{j}$ NWRP . So the number $n$ of nodes corresponds to the number of components in the composite system involved in the network, the number $r$ of factors contained in the product state corresponds to the number of resources. The number of nodes involved in the resource with the most shared nodes $k=\sharp(L)=\max\limits_{i}\sharp(L_{i})$ as the depth of the network. It reflects the fact that the quantum network state with respect to this network can be obtained by preparing the states in at most $k$-partite system, and the corresponding network states of the form in Eq.(6) are called network states with depth $k$.
Example 7.
Consider the scenario that we have three networks, net 1, net 2 and net 3, as in Fig.1, and three 6-qubit states,
$\rho_{a}=|\psi_{a}\rangle\langle\psi_{a}|,\ \ \rho_{b}=|\psi_{b}\rangle\langle\psi_{b}|,\ \ \rho_{c}=|\psi_{c}\rangle\langle\psi_{c}|$
determined by these networks, where
$\begin{array}[]{rl}|\psi_{a}\rangle=&\frac{1}{8}(|000000\rangle+|000011\rangle+|001100\rangle+|001111\rangle\\ &+|110000\rangle+|110011\rangle+|111100\rangle+|111111\rangle),\end{array}$
$|\psi_{b}\rangle=\frac{1}{4}(|000000\rangle+|000111\rangle+|111000\rangle+|111111\rangle),$
$|\psi_{c}\rangle=\frac{1}{4}(|000000\rangle+|000011\rangle+|111100\rangle+|111111\rangle).$
The task is that, for each $j\in\{1,2,3\}$, find $x\in\{a,b,c\}$ such that $\rho_{x}$ is exactly the network state corresponding to the network net $j$.
Figure 1: Three different quantum networks.
In this example, we also take $s=$-$\infty$ for an illustration. To do this, we observe that, for the network net 1 in Fig.1, parties $A,B$ and $C$ share a resource state $\rho_{1}$, while parties $C,D$ and $E$ share a resource state $\rho_{2}$. This indicates that the depth of the network net 1 is 3 and the corresponding network state must be 3-producible. Similarly, the depth of the network net 2 is 2 and the corresponding network state is 2-producible, while the depth of the network net 3 is 4 and the corresponding network state is 4-producible.
It is known by Theorem 4, for a state $\rho$ in a $6$-partite system and $1\leq k\leq 5$, $\rho$ is $k$-producible implies that
$\sum_{b=1}^{3}\sum_{n=1}^{2}I^{s}(\rho,{\bf P}_{6,n}^{(b)})\leq I_{k},$
where
$I_{k}=\begin{cases}6k(\kappa-\frac{1}{2})+6(2\kappa-1)&{\rm if}\ \ 6=pk;\\ (36+p^{2}k^{2}+pk^{2}-12pk)(\kappa-\frac{1}{2})+6(2\kappa-1)&{\rm if}\ \ 0<6-pk<k,\end{cases}$
$p=[\frac{6}{k}]$. The concrete values of $I_{k}$ for each $k$ is list in Table IX.
$k$ 1 2 3 4 5
$I_{k}$ 9 12 15 16 19
Table 9: Threshold values $I_{k}$ for $k$-producible states in 6-partite quantum systems.
Regard $\rho_{a}$, $\rho_{b}$ and $\rho_{c}$ as $6$-partite states. A direct calculation gives
$I_{a}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{s}(\rho_{a},{\bf P}_{6,n}^{(b)})=12,$
$I_{b}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{s}(\rho_{b},{\bf P}_{6,n}^{(b)})=15$
and
$I_{c}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{s}(\rho_{c},{\bf P}_{6,n}^{(b)})=16.$
Comparing with Table IX, we have
$I_{1}<I_{a}\leq I_{k}\quad\mbox{\rm for }k\leq 2;$
$I_{1}<I_{2}<I_{b}\leq I_{k}\quad\mbox{\rm for }k\geq 3$
and
$I_{1}<I_{2}<I_{3}<I_{c}\leq I_{k}\quad\mbox{\rm for }k\geq 4.$
Then, by Theorem 4, it is easily seen that $\rho_{c}$ is not 3-producible and thus is the network state corresponding to the network of depth greater than 3. Consequently, $\rho_{c}$ is the state corresponding to net 3. Similarly, one concludes that $\rho_{b}$ is the network state corresponding to the network net 1 and $\rho_{a}$ is corresponding to net 2. $\Box$
IV Conclusion
We establish two criteria based on MUMs and the generalized Wigner-Yanase skew information to detect $k$-nonseparability and $k$-partite entanglement, respectively, in multipartite quantum systems. Through two examples, we found that when the parameter takes the limiting value $s=$-$\infty$, the generalized Wigner-Yanase skew information exhibits the best detection performance, being more effective in detecting multipartite quantum entanglement compared to other values of $s$. These findings could offer new perspectives on detecting multipartite quantum entanglement and may also be applicable to the detection of other multipartite quantum correlations. It is also worth noting that the upper bounds of the inequalities obtained in our main theorems are independent of the parameter $s$ of the generalized Wigner-Yanase skew information $I^{s}(\rho,A)$. This is because we used the inequality $I^{s}(\rho,A)\leq V(\rho,A)$ during the proof process. However, given the superior performance at the limiting case $s=$-$\infty$, an important and challenging direction for future research is whether one can establish tighter upper bounds that depend on $s$, thereby further enhancing the detection capability of multipartite entanglement.
V Acknowledgments
This work is partially supported by the National Natural Science Foundation of China (12171290, 12071336).

### v3

@@
-
+Detecting $k$-nonseparability and $k$-partite Entanglement with Generalized Skew Information and Mutually Unbiased Measurements
 Abstract
@@
 The structure of this paper is as follows.
-In Section 2, we give some background knowledge on several quantum informations, $k$-separability, and $k$-partite entanglement in multipartite quantum systems.
+In Section 2, we give some background knowledge on quantum informations, $k$-separability, and $k$-partite entanglement in multipartite quantum systems.
 Section 3 is devoted to derive criteria for detecting $k$-nonseparability and $k$-partite entanglement in multipartite quantum systems by combining MUMs and the generalized Wigner-Yanase skew information.
@@
 Denote by $\mathcal{B}(H)$ the algebra of all bounded linear operators and ${\mathcal{S}}(H)=\{\rho:\rho\geq 0,{\rm Tr}(\rho)=1\}$ is the closed convex set of all quantum states in $\mathcal{B}(H)$.
-$\mathbb{I}\in{\mathcal{B}}(H)$ and $\mathbb{I}_{i}\in{\mathcal{B}}(H_{i})$ stand for the identity operator on $H$ and $H_{i}$, respectively.
+$\mathbb{I}\in{\mathcal{B}}(H)$ and $\mathbb{I}_{i}\in{\mathcal{B}}(H_{i})$ stand for the identity operators on $H$ and $H_{i}$, respectively.
 II.1 $k$-nonseparability and $k$-partite entanglement
@@
 where $\gamma_{1}|\gamma_{2}|\cdots|\gamma_{m}$ is a $m$-partition of $\{1,2,\cdots,N\}$, and the number of particles in each subset $\gamma_{i}$ is not greater than $k$.
-For any mixed state $\rho$, we say that $\rho$ is $k$-producible if it can be represented as a convex combination of $k$-producible pure states; otherwise, we say that $\rho$ contains $(k+1)$-partite entanglement GOTG .
+For any mixed state $\rho$, we say that $\rho$ is $k$-producible if it can be represented as a convex combination of $k$-producible pure states; otherwise, we say that $\rho$ contains $(k+1)$-partite entanglement, or is $(k+1)$-partite entangled GOTG .
 Denote by $\mathcal{S}_{k}$ the set of all $k$-separable quantum states ($2\leq k\leq N$) and by $\mathcal{P}_{k}$ the set of all $k$-producible quantum states ($1\leq k\leq N-1$).
@@
 II.5 Mutually unbiased measurements (MUMs)
-Let $P^{(b)}=\{P_{n}^{(b)}:n=1,\cdots,d\}$ be two positive-operator-valued measurements (POVMs), $b=1,2$. $P^{(1)}$ and $P^{(2)}$ are called MUMs if they satisfy the following conditions:
+Let $P^{(b)}=\{P_{n}^{(b)}:n=1,\cdots,d\}\ (b=1,2)$ be two positive-operator-valued measurements (POVMs) on a $d$-dimensional Hilbert space $H_{0}$.
+$P^{(1)}$ and $P^{(2)}$ are called MUMs if they satisfy the following conditions:
 $\begin{cases}\mathrm{tr}(P_{n}^{(b)})=1,\\ \mathrm{tr}(P_{n}^{(b)}P_{n^{\prime}}^{(b^{\prime})})=&\delta_{n,n^{\prime}}\delta_{b,b^{\prime}}\kappa\\ &+(1-\delta_{n,n^{\prime}})\delta_{b,b^{\prime}}\frac{1-\kappa}{d-1}+(1-\delta_{b,b^{\prime}})\frac{1}{d},\end{cases}$
@@
 Kalev and Gour in KG proposed a construction method for MUMs in any dimensional quantum system.
-Take an orthonormal basis $\{F_{k}\}_{k=1}^{d^{2}-1}$ consisting of Hermitian and traceless operators acting on $H$.
-Arrange these basis elements on a grid of d1d-1 columns and $d+1$ rows, and denote them as $\{F_{n,b}\}$ with $n=1,2,\ldots,d-1$ and $b=1,2,\ldots,d+1$.
-Define $d(d+1)$ operators $\{F_{n}^{(b)}\}$ by
+Take an orthonormal basis $\{F_{k}\}_{k=1}^{d^{2}-1}$ for $B_{0}(H_{0})$, consisting of Hermitian and traceless operators acting on $H_{0}$ and with inner product $(A,B)={\rm Tr}(B^{\dagger}A)$.
+Arrange these basis elements on a grid of d1d-1 columns and $d+1$ rows: the first row consists of $F_{1},F_{2},\ldots,F_{d-1}$, the second row consists of $F_{d},F_{d+1},\ldots,F_{2(d-1)}$, and so on.
+Denote the element in the $n$-th column of the $b$-th row by $F_{n,b}$, with $n=1,2,\ldots,d-1$ and $b=1,2,\ldots,d+1$.
+Define $d(d+1)$ operators $F_{n}^{(b)}$ by
 $F_{n}^{(b)}=\left\{\begin{array}[]{ll}F^{(b)}-(d+\sqrt{d})F_{n,b}&\text{for}\ n=1,\cdots,d-1,\\ (1+\sqrt{d})F^{(b)}&\text{for}\ n=d,\end{array}\right$. (1)
@@
 $\sum_{n=1}^{d}F_{n}^{(b)}=0,\ \ \forall b\in\{1,2,\cdots,d+1\}.$ (2)
-The measurement operators $\{P_{n}^{(b)}\}$ are given by:
-$P_{n}^{(b)}=\frac{1}{d}\mathbb{I}+tF_{n}^{(b)},$ (3)
-where the free parameter $t$ is chosen such that $P_{n}^{(b)}\geq 0$; and
-$\kappa=\frac{1}{d}+t^{2}(1+\sqrt{d})^{2}(d-1).$ (4)
-Then $\{P^{(b)}=\{P_{n}^{(b)}\}\}$ form a complete set of $d+1$ mutually unbiased measurements, where $b$ labels the measurement and $n$ labels the outcome.
-Moreover, any complete set of MUMs has this form.
-It was also proven in LL that $\left\{\frac{\mathbb{I}}{\sqrt{d}},F_{n,b}:n=1,\cdots,d-1;b=1,\cdots,d+1\right\}$ forms a set of local orthogonal observables (LOOs) on $H$; and any complete set of MUMs $\{P_{n}^{(b)}:n=1,2,\cdots,d;b=1,2,\cdots,d+1\}$ must satisfy the following equation:
+It was proved in KG that there exists a parameter $t$, called the free parameter, such that the operators $P_{n}^{(b)}$ given by
+$P_{n}^{(b)}=\frac{1}{d}\mathbb{I}+tF_{n}^{(b)}$ (3)
+are positive for all $n,b$. Put
+$\kappa=\frac{1}{d}+t^{2}(1+\sqrt{d})^{2}(d-1),$ (4)
+then $P^{(b)}=\{P_{n}^{(b)}\}_{n=1}^{d}$ ($b=1,2,\cdots,d+1$) form a complete set of $d+1$ MUMs, where $b$ labels the measurement and $n$ labels the outcome.
+Moreover, any complete set of MUMs has this form and satisfies the following equation:
 $\sum_{b=1}^{d+1}\sum_{n=1}^{d}(P_{n}^{(b)})^{2}=(d+1)\kappa\mathbb{I}.$ (5)
+It was also shown in $L$ that
+$\left\{\frac{\mathbb{I}}{\sqrt{d}},F_{n,b}\colon n=1,\dots,d-1;\,b=1,\dots,d+1\right\}$
+forms a set of local orthogonal observables (LOOs) on $H_{0}$.
 III $k$-separability and $k$-partite entanglement detection criteria
 In this section, we will establish some entanglement detection criteria for $k$-separability and $k$-partite entanglement in multipartite quantum systems.
-For convenience, assume $H_{i}=H_{0}$ with $\dim H_{0}=d$, $i=1,2,\cdots,N$. Take the same set of $d+1$ MUMs acting on each subsystem $H_{i}$
-$\{P^{(i,b)}=\{P_{i,n}^{(b)}\}:b=1,2,\cdots,d+1;n=1,2,\cdots,d\},$
+For convenience, assume $H_{i}=H_{0}$ with $\dim H_{0}=d$, $i=1,2,\dots,N$. Take the same $d+1$ MUMs on each subsystem $H_{i}$:
+$P^{(i,b)}=\{P_{i,n}^{(b)}\}_{n=1}^{d}\ \ (b=1,2,\cdots,d+1),$
 where $i$ is used to label the subsystem on which the operators act, $i=1,2,\cdots,N$.
@@
 $=$ $\frac{d+1}{d}(\otimes_{i\in\Gamma}\mathbb{I}_{i})+t^{2}(d+\sqrt{d})^{2}\sum_{b=1}^{d+1}\sum_{n=1}^{d-1}(\otimes_{j=1}^{p-1}\mathbb{I}_{j})\otimes F_{n,b}\otimes(\otimes_{j=p+1}^{q-1}\mathbb{I}_{j})\otimes F_{n,b}\otimes(\otimes_{j=q+1}^{\sharp(\Gamma)}\mathbb{I}_{j})$
-$\leq$ $\frac{d+1}{d}(\otimes_{i\in\Gamma}\mathbb{I}_{i})+t^{2}(d+\sqrt{d})^{2}(1-\frac{1}{d})(\otimes_{i\in\Gamma}\mathbb{I}_{i})=(1+\kappa)(\otimes_{i\in\Gamma}\mathbb{I}_{i}).$
-$\Box$
+$\leq$ $\frac{d+1}{d}(\otimes_{i\in\Gamma}\mathbb{I}_{i})+t^{2}(d+\sqrt{d})^{2}(1-\frac{1}{d})(\otimes_{i\in\Gamma}\mathbb{I}_{i})$
+$=$ $(1+\kappa)(\otimes_{i\in\Gamma}\mathbb{I}_{i}).$
+Here, the first equality uses Eq.(3), the third equality uses Eq.(2), the fourth equality uses Eq.(1) and the last equality uses Eq.(4). $\Box$
 Proposition 2.
@@
 $g_{s}(\rho)=\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}\left(\rho,\mathbf{P}_{N,n}^{(b)}\right),\quad s\in[-\infty,0].$
-Since $\rho$ is $k$-separable, then $\rho$ can be written as a convex combination of $k$-separable pure states, that is, $\rho=\sum_{r}p_{r}\rho_{r}$, where $\rho^{(r)}$ has the form
+Since $\rho$ is $k$-separable, $\rho$ can be written as a convex combination of $k$-separable pure states, that is, $\rho=\sum_{r}p_{r}\rho_{r}$, where $\rho^{(r)}$ has the form
 $\rho_{r}=\rho^{(r)}_{\Gamma_{1}}\otimes\rho^{(r)}_{\Gamma_{2}}\otimes\cdots\otimes\rho^{(r)}_{\Gamma_{k}}$
@@
 It is obvious that $y_{0}\in X$.
-Recall that, for any vectors $x=(x_{1},x_{2},\cdots,x_{k})^{t}$ and $y=(y_{1},y_{2},\cdots,y_{k})^{k}\in{\mathbb{R}}^{k}$, $x\prec y$ means that $\sum_{i=1}^{m}x_{i}\leq\sum_{i=1}^{m}y_{i}$ holds for all $m=1,2,\cdots,k-1$ and $\sum_{i=1}^{k}x_{i}=\sum_{i=1}^{k}y_{i}$.
+Recall that, for any nonnegative vectors $x=(x_{1},x_{2},\cdots,x_{k})^{t}$ and $y=(y_{1},y_{2},\cdots,y_{k})^{t}\in{\mathbb{R}}^{k}$ with coordinate descending order arrangement, $x\prec y$ means that $\sum_{i=1}^{m}x_{i}\leq\sum_{i=1}^{m}y_{i}$ holds for all $m=1,2,\cdots,k-1$ and $\sum_{i=1}^{k}x_{i}=\sum_{i=1}^{k}y_{i}$.
 Then we have $x\prec y_{0}$ for all $x\in X$; otherwise, if there exists some $m^{\prime}$ such that $\sum_{i=1}^{m^{\prime}}x_{i}>\sum_{i=1}^{m^{\prime}}y_{i}=N-k+m^{\prime},$ then there must be some $x_{0}=0$, a contradiction.
@@
 Thus, if $N=pk$, then
-$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ &\leq pk^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)=Nk(\kappa-\frac{1}{d})+N(d\kappa-1);\end{array}$
+$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ \leq&pk^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)=Nk(\kappa-\frac{1}{d})+N(d\kappa-1);\end{array}$
 and if $N=pk+m$ with $0<m<k$, then
-$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ &\leq pk^{2}(\kappa-\frac{1}{d})+(N-pk)^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)\\ &=(N^{2}+p^{2}k^{2}+pk^{2}-2pkN)(\kappa-\frac{1}{d})+N(d\kappa-1).\end{array}$
+$\begin{array}[]{rl}&\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})\\ \leq&pk^{2}(\kappa-\frac{1}{d})+(N-pk)^{2}(\kappa-\frac{1}{d})+N(d\kappa-1)\\ =&(N^{2}+p^{2}k^{2}+pk^{2}-2pkN)(\kappa-\frac{1}{d})+N(d\kappa-1).\end{array}$
 Thus, the monotonicity of generalized Wigner-Yanase skew information ensures that the tightest bounds are achieved in the limiting case $s=$-$\infty$. This completes the proof of the theorem. $\Box$
 Consequently, any violation of the inequalities in Theorems 3-4 indicate $k$-nonseparability and $(k+1)$-partite entanglement, respectively.
+We remark here that, although the supremum over $s\in[-\infty,0]$ in Theorems 3-4 is theoretically difficult to evaluate due to the nonlinear dependence of $\sum_{b=1}^{d+1}\sum_{n=1}^{d}I^{s}(\rho,{\bf P}_{N,n}^{(b)})$ on $s$, in practice it can be well approximated by considering extreme values such as $s=$-$\infty$ or by sampling $s$ at a finite set of points.
+This allows effective detection of $k$-nonseparability while retaining the practical advantages of the criterion.
 To verify the validity of these criteria, we provide several examples below.
 Example 5.
-Consider the 6-partite quantum state
+Consider the 6-partite quantum state on the Hilbert space $H=(\mathbb{C}^{2})^{\otimes 6}$:
 $\rho(p)=p|\psi\rangle\langle\psi|+\frac{1-p}{64}\mathbb{I},$
-where $|\psi\rangle=\frac{1}{\sqrt{6}}(|000001\rangle+|000010\rangle+|000100\rangle+|001000\rangle+|010000\rangle+|100000\rangle)$.
+where $|\psi\rangle=\frac{1}{\sqrt{6}}(|000001\rangle+|000010\rangle+|000100\rangle+|001000\rangle+|010000\rangle+|100000\rangle)$ and $0\leq p\leq 1$.
 In this case, we illustrate the situation by taking s=1,0,s=-1,0,-$\infty$, respectively.
@@
 Example 6.
-Consider an 11-partite quantum state
-$\rho=p|\psi\rangle\langle\psi|+\frac{1-p}{2^{11}}\mathbb{I}$
-where $|\psi\rangle=\frac{1}{\sqrt{2}}(|0\rangle^{\otimes 11}+|1\rangle^{\otimes 11})$.
+Consider an 11-partite quantum state on the Hilbert space $H=(\mathbb{C}^{2})^{\otimes 11}$:
+$\rho=p|\psi\rangle\langle\psi|+\frac{1-p}{2^{11}}\mathbb{I},$
+where $|\psi\rangle=\frac{1}{\sqrt{2}}(|0\rangle^{\otimes 11}+|1\rangle^{\otimes 11})$ and $0\leq p\leq 1$.
 In this example, we take $s=$-$\infty$ for an illustration.
 According to Theorem 3 and Theorem 4, we can conclude that $\rho(p)$ is $k$-nonseparable if $p_{k}<p\leq 1$ and exhibits $(k+1)$-partite entanglement if $\tilde{p_{k}}<p\leq 1$.
-Note that, $p_{k}^{\prime}$ is the result obtained using the method in HLS , and $\tilde{p_{k}^{\prime}}$ is obtained in HQGY , as shown in the Tables VII-VIII.
+Note that, $p_{k}^{\prime}$ is the result obtained using the method in HLS , and $\tilde{p_{k}^{\prime}}$ is obtained in HQGY2 , as shown in the Tables VII-VIII.
 $k$ 2 3 4 5 6 7 8 9 10 11
@@
 Generally, a quantum network consists of $n$ nodes (subsystems, denote by $A_{j}$, $j=1,2,\cdots,n$) and $r$ quantum resources $\rho_{i}$ shared by some of the nodes.
-Let $L_{i}\subseteq\{1,2,\cdots,n\}$ the subset of index of the nodes that $\{A_{j}:j\in L_{i}\}$ is exact the subset of all nodes that share the $i$-th resource $\rho_{i}$.
+Let $L_{i}\subseteq\{1,2,\cdots,n\}$ be the subset of index of the nodes that $\{A_{j}:j\in L_{i}\}$ is exact the subset of all nodes that share the $i$-th resource $\rho_{i}$.
 Denote by $\rho_{i}=\rho_{L_{i}}$, $i=1,2,\cdots,r$.
 Then the corresponding quantum network state has the form:
-$\rho=\rho_{L_{1}}\otimes\rho_{L_{2}}\otimes\cdots\rho_{L_{r}}.$
+$\rho=\rho_{L_{1}}\otimes\rho_{L_{2}}\otimes\cdots\otimes\rho_{L_{r}}.$
 Accordingly, for every $j=1,2,\ldots,n$, the node $A_{j}$ may share several $\rho_{L_{i}}$ with other nodes, that is $j\in L_{i}$.
 Let $F_{j}=\{i_{j}:A_{j}\in L_{i}\}\subseteq\{1,2,\ldots,r\}$.
-Then the subsystem hold by $A_{j}$ is a $\sharp F_{j}$-partite composite system $H_{A_{j}}=\bigotimes_{i_{j}\in F_{j}}H_{A_{j_{i}}}$.
+Then the subsystem hold by $A_{j}$ is a $\sharp F_{j}$-partite composite system $H_{A_{j}}=\bigotimes_{i_{j}\in F_{j}}H_{A_{i_{j}}}$.
 Every quantum network state determined by this network has the form:
-$\rho=\sum_{\lambda\in\Lambda}p_{\Lambda}(\lambda)(\Omega_{1}^{\lambda}\otimes\cdots\otimes\Omega_{n}^{\lambda})(\rho_{L_{1}}^{\lambda}\otimes\rho_{L_{2}}^{\lambda}\otimes\cdots\rho_{L_{r}}^{\lambda}),$ (6)
+$\rho=\sum_{\lambda\in\Lambda}p_{\Lambda}(\lambda)(\Omega_{1}^{\lambda}\otimes\cdots\otimes\Omega_{n}^{\lambda})(\rho_{L_{1}}^{\lambda}\otimes\rho_{L_{2}}^{\lambda}\otimes\cdots\otimes\rho_{L_{r}}^{\lambda}),$ (6)
 where $\{p_{\Lambda}(\lambda)\}$ is a probability distribution, $\Omega_{j}^{\lambda}$ is a quantum channel of the subsystem $A_{j}$ NWRP .
-So the number $n$ of nodes corresponds to the number of components in the composite system involved in the network, the number $r$ of factors contained in the product state corresponds to the number of resources.
-The number of nodes involved in the resource with the most shared nodes $k=\sharp(L)=\max\limits_{i}\sharp(L_{i})$ as the depth of the network.
-It reflects the fact that the quantum network state with respect to this network can be obtained by preparing the states in at most $k$-partite system, and the corresponding network states of the form in Eq.(6) are called network states with depth $k$.
+So the number $n$ of nodes corresponds to the number of components in the composite system involved in the network, the number $r$ of factors contained in the product state corresponds to the number of sources.
+Denote by $\sharp(L_{i})$ the number of nodes involved in the source $\rho_{i}$.
+The most shared nodes $k=\max\limits_{i}\sharp(L_{i})$ is called the depth of the network, which reflects the fact that the quantum network state with respect to this network can be obtained by preparing the states in at most $k$-partite system, and the corresponding network states of the form in Eq.(6) are called network states with depth $k$.
 Example 7.
@@
 In this example, we also take $s=$-$\infty$ for an illustration.
-To do this, we observe that, for the network net 1 in Fig.1, parties $A,B$ and $C$ share a resource state $\rho_{1}$, while parties $C,D$ and $E$ share a resource state $\rho_{2}$.
+To do this, we observe that, for the network net 1 in Fig.1, parties $A,B$ and $C$ share a source state $\rho_{1}$, while parties $C,D$ and $E$ share a source state $\rho_{2}$.
 This indicates that the depth of the network net 1 is 3 and the corresponding network state must be 3-producible.
@@
 Regard $\rho_{a}$, $\rho_{b}$ and $\rho_{c}$ as $6$-partite states. A direct calculation gives
-$I_{a}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{s}(\rho_{a},{\bf P}_{6,n}^{(b)})=12,$
-$I_{b}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{s}(\rho_{b},{\bf P}_{6,n}^{(b)})=15$
+$I_{a}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{-\infty}(\rho_{a},{\bf P}_{6,n}^{(b)})=12,$
+$I_{b}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{-\infty}(\rho_{b},{\bf P}_{6,n}^{(b)})=15$
 and
-$I_{c}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{s}(\rho_{c},{\bf P}_{6,n}^{(b)})=16.$
+$I_{c}=\sum_{b=1}^{3}\sum_{n=1}^{2}I^{-\infty}(\rho_{c},{\bf P}_{6,n}^{(b)})=16.$
 Comparing with Table IX, we have
@@
 Similarly, one concludes that $\rho_{b}$ is the network state corresponding to the network net 1 and $\rho_{a}$ is corresponding to net 2.
+This completes the given task.
 $\Box$
@@
 V Acknowledgments
-This work is partially supported by the National Natural Science Foundation of China (12171290, 12071336).
+The authors wish to give their thanks to the referees for many helpful comments to improve the original paper.
+This work is partially supported by the National Natural Science Foundation of China (12571338, 12171290, 12071336).